### v1

# Entanglement-breaking channels are a quantum memory resource

Abstract
Entanglement-breaking channels (equivalently, measure-and-prepare channels) are an important class of quantum operations noted for their ability to destroy multipartite spatial quantum correlations. Inspired by this property, they have also been employed in defining notions of “classical memory”, under the assumption that such channels effectively act as a classical resource. We show that, in a single-system multi-time scenario, entanglement-breaking channels are still a quantum memory resource: a qudit going through an entanglement-breaking channel cannot be simulated by a classical system of same dimension. We provide explicit examples of memory-based output generation tasks where entanglement-breaking channels outperform classical memories of the same size. Our results imply that entanglement-breaking channels cannot be generally employed to characterize classical memory effects in temporal scenarios without additional assumptions.
Introduction.— The notion of classical memory, in the sense of storage of information, is familiar to us from our everyday life. When we talk about the number of bits (or, more commonly, one of its multiples: bytes, kilobytes, etc.) a device has, we refer the logarithm of the number of distinguishable internal states, where information can be stored and retrieved. Needless to say, memory is an essential resource for information processing, and a finite amount of memory imposes fundamental limitations on which tasks can be achieved. With the advent of quantum information and quantum computation, distinguishing between classical and quantum memory effects is of utmost importance for identifying scenarios where quantum systems offer an advantage over classical ones [1, 2, 3, 4]. An approach towards this goal is by establishing bounds on information processing tasks achievable with classical memory alone. A violation of such bounds then provides a certificate of non-classicality, often also providing a quantifiable metric for the quantumness of the underlying memory resource [5, 6, 7, 8, 9, 10, 11, 12].
Any information processing task consists of manipulating information stored in a memory over time. This is achieved by sequential operations performed on the memory, with classical and quantum systems offering vastly different capabilities on what memory states and their transformations are possible. In recent years, the formalism of quantum supermaps [13, 14, 15], also known as process tensors or quantum combs, has enabled a thorough investigation of temporal quantum protocols, as it is naturally well-suited for the study of multi-time statistics [7, 16, 17]. Within this framework, many quantum-inspired notions have been employed as conflicting definitions of “classical memory” [1, 16, 17, 4, 18]; see Ref. [17] for a detailed characterization.
A typical approach involves entanglement-breaking (EB) channels, which are known to be exactly equivalent to measure-and-prepare (M&P) operations [19]. As the input quantum state is destroyed by the measurement operation and replaced by a classical label (albeit encoded in the prepared output quantum state), one may be inclined to claim that such channels are essentially equivalent to classical channels [1, 9, 10, 11, 16], i.e., that they can only establish classical correlations between their inputs and outputs. EB channels have been used in this manner, e.g., to develop resource theories of quantum memories [1], the benchmarking and application of quantum devices [9], characterizing non-classical memory effects in quantum processes [4, 20, 16] and bosonic Gaussian channels [21, 11], and in establishing a hierarchy between competing notions of classical memory [17]. A potential issue with such entanglement-inspired definitions is that they still operate within the quantum formalism, and therefore, may not truly characterize the limitations inherent to classical systems. Clearly, a precise distinction between classical and quantum memory effects must emerge from explicit constructions within each theory.
Here, we show that the previous interpretation and applications of EB channels can be misleading in the case of temporal scenarios involving sequential measurements. We address the following question: is the memory of a “$d$-it” (a classical system with $d$ states) equivalent to that of a qudit ($d$-level quantum system) passing through an EB channel? We answer it in the negative by presenting explicit examples where a qudit passing through an EB channel outperforms a $d$-it of classical memory in the task of generating correlations in time. We conclude that EB channels cannot be used to define or characterize classical memories in the most general settings.
Figure 1: Temporal correlations can be investigated by a classical or quantum system with $d$ distinguishable internal states, in the initial state $s_{0}$, which is repeatedly measured by the same instrument at multiple times. The measurement of state $s_{t}$ produces the outcome $a_{t+1}$ and changes the state to $s_{t+1}$.
Preliminary notions.— A channel $\mathcal{E}_{A}$ is said to be entanglement breaking if $(\mathcal{E}_{A}\otimes\text{id}_{B})(\rho_{AB})$ is separable for any bipartite state $\rho_{AB}$. A well-known result establishes that any EB channel $\mathcal{E}$ can be written as a measure-and-prepare (M&P) channel in the Holevo form [22, 19]
$\mathcal{E}(\rho)=\sum_{i=0}^{m}\operatorname{tr}\left[\rho E_{i}\right]\sigma_{i},$ (1)
for $\{E_{i}\}_{i=1}^{m}$ a Positive Operator-Valued Measure (POVM), i.e. $E_{i}\geq 0$ and $\sum_{i=1}^{m}E_{i}=\mathbbm{1}$, and $\sigma_{i}$ quantum states. Here, $m$ can be interpreted as the number of labels—in principle, arbitrarily large—required in the channel’s intermediate classical memory in order to describe its operation [19, 1]. EB channels can be fully characterized [19] by the fact their Choi matrix [23, 24], $C_{\mathcal{E}}:=(\mathcal{E}\otimes\text{id})(\left|\Phi\right\rangle\left\langle\Phi\right|)$ with $\left|\Phi\right\rangle$ a maximally entangled state, is separable and with rank $r\geq d$, which also implies any such channel can be rewritten with at most $m=d^{2}$ operators $E_{i}$ and $\sigma_{i}$. These channels can be understood in terms of a quantum-classical-quantum operation: (1) The input quantum state $\rho$ is first measured, producing a classical outcome label $i$, out of a total $m$, with probability $\operatorname{tr}\left[\rho E_{i}\right]$, (2) Conditioned on $i$, the corresponding quantum state $\sigma_{i}$ is passed forward as output; see Fig. 2b.
To properly distinguish between classical and quantum memories, we investigate their capabilities when generating temporal correlations in the simplest scenario, i.e., that of a single, finite-dimensional system subject to the same dichotomous measurement at multiple times, as shown in Fig. 1. Such a system can be described within the formalism of finite-state machines [6, 25]. In the classical case, the most general description of its behavior is by an initial stochastic vector $\pi$ over its $d$ distinct states (i.e., a $d$-it of memory), and a pair of (row) sub-stochastic transition matrices $T=(T_{0},T_{1})$ describing its dynamics upon each outcome $a\in\{0,1\}$, with $T_{0}+T_{1}$ row stochastic. The probability of a length-$L$ sequence of outcomes $\bm{a}=a_{1}a_{2}\cdots a_{L}$ is then given by
$p(\bm{a}|T,d)=\pi T_{a_{1}}\cdots T_{a_{L}}\eta,$ (2)
with $\eta=(1,\dots,1)^{\text{T}}$. The quantum case is defined analogously, with a single $d$-dimensional quantum system (a qudit), initially at state $\rho_{0}$, being repeatedly measured by a quantum instrument $\mathcal{I}=(\mathcal{I}_{0},\mathcal{I}_{1})$, with effects $\mathcal{I}_{a}$ being completely positive (CP) trace non-increasing maps such that $\mathcal{I}_{0}+\mathcal{I}_{1}$ is completely positive trace-preserving (CPTP). The probability of a sequence $\bm{a}$ is given by
$p(\bm{a}|\mathcal{I},d)=\operatorname{tr}\left[\mathcal{I}_{a_{L}}\circ\cdots\circ\mathcal{I}_{a_{1}}(\rho_{0})\right].$ (3)
In both cases, the system’s state is the sole resource available to generate the probability distributions, and any dynamics is entirely governed by the effects associated with an outcome $a_{t}$. In this way, classical and quantum memories are placed on equal footing as a memory resource. Crucially, memory is the central resource for this task, as any arbitrary temporal correlation can be generated by a classical system with enough memory, i.e., enough internal states [5, 26, 27]. Consequently, the system’s dimension establishes fundamental limitations on the set of achievable temporal correlations [28, 6, 27, 8, 25], providing a clear notion for distinguishing between manifestly classical and quantum memories. Note that we do not allow for external storage of information (e.g., by conditioning on past outcomes, or invoking an external reference clock), so that the total amount of memory can be quantified by the number of states $d$ of the system.
One possible approach to distinguish between classical and quantum memory effects is by establishing upper bounds on the maximum probability for individual sequences to be produced within classical and quantum mechanics [6, 8, 25]. The notion of deterministic complexity of a sequence $\bm{a}$, denoted by $\operatorname{DC}(\bm{a})$, was introduced in Ref. [25] as the minimum dimension $d$ for a system (classical or quantum) to be capable of generating $\bm{a}$ with probability one. Thus, if $d<\operatorname{DC}(\bm{a})$, then no classical or quantum system can produce the sequence $\bm{a}$ deterministically, implying both classical and quantum theory must obey non-trivial upper bounds for this maximum probability. Since $d=2$ is the smallest non-trivial amount of memory, these observations also highlight the importance of studying temporal correlations over three or more time steps, as every sequence of length two can be produced deterministically with either a bit or qubit.
As the underlying memory resource can be explicitly confined within each theory, such non-trivial upper bounds provide a sharp demarcation between classical and quantum memories and their capabilities, but they are generally difficult to obtain precisely [25, 29, 30]. In Ref. [25], a universal upper bound of $1/\mathrm{e}$ was conjectured for all sub-deterministic classical memory scenarios, whereas no non-trivial universal quantum bound seems to exist.
Figure 2: (a) The sequential measurement protocol considered in this work. An isolated quantum system in an initial state $\rho_{0}$ is repeatedly passed through an entanglement-breaking channel $\mathcal{E}$ before being measured by a quantum instrument $\mathcal{I}$, obtaining a sequence of outcomes $\bm{a}=a_{1}a_{2}\dots a_{L}$. (b) The entanglement-breaking channel $\mathcal{E}$ understood as a measure-and-prepare operation on a $d$-dimensional quantum system, with a $m$-dimensional intermediate classical space. (c) The $m\times m$ transition matrix of the effective classical model.
The sequential measurement protocol.— We investigate the non-classicality of entanglement-breaking channels by inserting an EB channel $\mathcal{E}$, as in Eq. 1, before each quantum measurement in Eq. 3 (depicted in Fig. 2a), obtaining
$\begin{split}&p(\bm{a}|\mathcal{I},\mathcal{E},d)=\operatorname{tr}\left[\mathcal{I}_{a_{L}}\circ\mathcal{E}\circ\cdots\circ\mathcal{I}_{a_{2}}\circ\mathcal{E}\circ\mathcal{I}_{a_{1}}\circ\mathcal{E}(\rho_{0})\right]\\ &=\sum_{i_{1}\dots i_{L}=1}^{m}\operatorname{tr}\left[\rho_{0}E_{i_{1}}\right]\operatorname{tr}\left[\mathcal{I}_{a_{1}}(\sigma_{i_{1}})E_{i_{2}}\right]\cdots\operatorname{tr}\left[\mathcal{I}_{a_{L}}(\sigma_{i_{L}})\mathbbm{1}\right]\\ &=\sum_{i_{1}\dots i_{L+1}=1}^{m}\pi_{i_{1}}[T_{a_{1}}]_{i_{1}i_{2}}\cdots[T_{a_{L}}]_{i_{L}i_{L+1}}\eta_{i_{L+1}},\end{split}$ (4)
where we use $[T_{a}]_{ij}:=\operatorname{tr}\left[\mathcal{I}_{a}(\sigma_{i})E_{j}\right]$, $\pi_{i}:=\operatorname{tr}\left[\rho_{0}E_{i}\right]$, and $\eta=(1,\dots,1)^{\text{T}}$ as in Eq. 2. Note that, since the $E_{i}$ are POVM elements, each trace term denoted a probability. We have thus collected all such probabilities in the matrices $T_{a}$ such that
$\sum_{a,j}[T_{a}]_{ij}=\operatorname{tr}\left[\textstyle{\sum_{a}}\mathcal{I}_{a}(\sigma_{i})\cdot\textstyle{\sum_{j}}E_{j}\right]=\operatorname{tr}\left[\hat{\sigma}_{i}\mathbbm{1}\right]=1,$ (5)
where we use the fact $E_{i}$ form a POVM and $\sum_{a}\mathcal{I}_{a}$ is CPTP, thus $\hat{\sigma}_{i}=\sum_{a}\mathcal{I}_{a}(\sigma_{i})$ has unit trace. Furthermore, we can define an initial probability vector $\pi$ arising from the first application of the channel, with $\sum_{i}\pi_{i}=1$. The above results imply we can model the protocol classically, by recognizing $T_{a}$ as $m\times m$ sub-stochastic transition matrices and $\pi$ as an initial stochastic vector (Fig. 2c). Eq. 4 can thus be equivalently written as in Eq. 2, with $m$ instead of $d$ states.
We are led to conclude that the repeated inclusion of an arbitrary entanglement-breaking channel before the quantum instrument allows the resulting temporal correlations to be, indeed, described classically. This observation seems to justify the association of EB channels with classical memories. Nevertheless, as we already discussed, any temporal correlation can be obtained via a classical system if enough memory is available, such that a meaningful comparison must be made between systems of bounded memory, i.e., bounded dimension.
A central observation is that the effective classical model thus obtained is over $m$ “virtual” classical states (i.e., the number of outcomes in the measure-and-prepare description of the channel), not $d$ as in the dimension of the underlying Hilbert space. This suggests that a qudit passing through an EB channel may be able to generate stronger correlations, i.e., carry more memory, than a classical $d$-it. Clearly, the exception is the case $m\leq d$, where any quantum model can be converted into a classical one of the same dimension (and vice versa) by choosing $E_{i},\sigma_{i}$, and $\mathcal{I}$ acting diagonally on the same basis; see App. A for the explicit construction.
Figure 3: (Color online). Probabilities of one-tick sequences $\bm{a}^{L}_{\text{ot}}$ for classical one-way models and the quantum construction using ETFs, both using $d=L-1$. One-way models are conjectured to be optimal for all $L=d-1$, and are known to achieve the exact upper bounds $\Omega_{C}(\bm{a}^{L}_{\text{ot}},d)$ for $L=3,4$, and $5$. The bounds for $L=4$ and $5$ are violated by this explicit quantum construction using ETFs, revealing non-classical memory effects of EB channels.
L d Classical ($\Omega_{C}$) Quantum ($\Omega_{Q}$) $\Omega_{Q}/\Omega_{C}$
3 2 $(2/3)^{3}=0.\overline{296}$ =?0.296296?${=}0.296296$ =?1?${=}1$
4 3 $(3/4)^{4}=0.31640625$ $\geq 0.359523$ $\geq 1.136270$
5 4 $(4/5)^{5}=0.32768$ $\geq 0.368445$ $\geq 1.124405$
Table 1: Results from numerical optimization for various one-tick sequences of length $L$, and Hilbert space dimension $d=L-1$. The quantum models found numerically for $L=4$ and $5$ outperform the previous ETF construction, and are included in App. D. No violations were found for $d=2$.
Quantum advantages with EB channels.— Here, we show that EB channels with $m>d$ are able to generate non-classical memory effects. This can be investigated by choosing a sequence $\bm{a}$ with $d<\operatorname{DC}(\bm{a})$, then computing the maximum probability achievable over all classical models with $d$ states:
$\Omega_{C}(\bm{a},d)$ $=\max_{T}\;p(\bm{a}|T,d).$ (6)
Analogously, using a $d$-dimensional quantum system, we look for instruments $\mathcal{I}$ and EB channels $\mathcal{E}$ with $m>d$ such that
$p(\bm{a}|\mathcal{I},\mathcal{E},d)>\Omega_{C}(\bm{a},d).$ (7)
Crucially, the maxima $\Omega_{C}(\bm{a},d)$ must be found explicitly within classical theory, either by exact computation or by obtaining an upper bound, whereas for the quantum value of $p(\bm{a}|\mathcal{I},\mathcal{E},d)$ an explicit model $(\mathcal{I},\mathcal{E})$ must be found. If the classical memory bound has been violated even in the presence of an EB channel $\mathcal{E}$, as in Eq. 7, then the system still acts as a quantum memory resource for the task of generating temporal correlations.
To achieve this goal, we consider the one-tick sequences [8, 25] of length $L$, defined as $\bm{a}_{\text{ot}}^{L}:=0\dots 01$ and having $\operatorname{DC}(\bm{a}^{L}_{\text{ot}})=L$, under the scenario $d=L-1$. As the case of $L=2$ is trivial, we consider the smallest non-trivial cases $L\geq 3$. For the classical model, analytical upper bounds can be computed via semidefinite programming (SDP) [29]. The case $L=3$ and $4$ were already solved in [29], and here we computed explicitly the case $L=5$ with the same method. In all these cases, the upper bounds computed via SDP coincide, up to the numerical precision, with the values obtained via the one-way model [8], which is conjectured to be optimal for all $d=L-1$ scenarios [8, 25]. We are therefore sure to have obtained the exact bounds for the cases $L=3,4,$ and $5$. An explicit quantum model violating these classical bounds can be obtained with the following construction. Consider the deterministic classical model on $m$ states given by $[T_{0}]_{i,i+1}=1$ for $i=1,\dots,m-1$, $[T_{1}]_{m,1}=1$, zeros for all other entries, and an initial state $\pi=(1,0,\dots,0)$. The idea is to approximate this model by emulating $m=d+1$ nearly-orthogonal virtual states with the available $d$-dimensional quantum system.
With $\zeta=\mathrm{e}^{2\pi\mathrm{i}/(m-1)}$, we define $\mathcal{E}$ using
$\left|\psi_{n}\right\rangle=\frac{1}{\sqrt{d-1}}\sum_{k=1}^{d-1}\zeta^{nk}\left|k\right\rangle,\quad n=1,\dots,m-1,$ (8)
then let $\sigma_{n}=\left|\psi_{n}\right\rangle\left\langle\psi_{n}\right|$ and $E_{n+1}=\tfrac{d-1}{d}\sigma_{n}$ for $n=1,\dots,m-1$, and finally setting $\rho_{0}=\sigma_{m}=E_{1}=\left|0\right\rangle\left\langle 0\right|$. The instrument is defined as $\mathcal{I}_{a}(\rho)=K_{a}\rho K^{\dagger}_{a}$ with $K_{0}=\mathrm{diag}(1,\cdots,1,0)$ and $K_{1}=\mathbbm{1}-K_{0}$. This quantum model corresponds to isolating the $(d-1)$-dimensional subspace orthogonal to $\left|0\right\rangle$, assembling a harmonic complex equiangular tight frame (ETF) [31] of size $m-1$ within this subspace, then using it for the virtual states in the transitions of $T_{0}$. ETFs are useful for this task as they provide a generalization of orthonormal bases, saturating the Welch bounds for minimum overlaps [32]. As seen in Fig. 3, this quantum ETF construction clearly violates the exact classical bounds for $L=4$ and $5$, conclusively showing that general EB channels cannot be applied to characterize classical memory effects.
The quantum model presented above provides an explicit violation of the classical bound. This is, however, not necessarily the maximum one can achieve in the quantum case. To investigate the quantum maximum with EB channels, we have also performed numerical optimization via gradient descent methods, obtaining lower bounds for the maxima $\Omega_{Q}(\bm{a},d)$. Results are summarized in Table 1. The quantum models found numerically slightly outperform the previous ETF construction. See App. C for more details regarding the optimization and App. D for the explicit models found.
Conclusion and discussion.— Entanglement breaking channels are universally recognized as classical memories [1, 2, 3, 9, 16, 4, 10, 17, 12, 11], and, indeed they are classical under many points of view. Quite surprisingly, however, we found that there exist scenarios where they still act as genuine quantum memories. More precisely, we answer in the negative the following question: is a qudit passing through an entanglement-breaking channel equivalent to a classical $d$-it?
To address this question, we have investigated a task where memory is the central resource, namely, the generation of correlations in time. Indeed, it has been shown that any temporal correlation can be reproduced by a classical system if enough memory is available [5, 26, 27], such that differences between classical and quantum temporal correlations only emerge if memory is bounded [6, 8, 25]. As classical memory is understood in terms of the number of internal states $d$ of the system, a fair comparison between classical and quantum theory must involve a quantum system of the same size. Differences between quantum and classical memories can be found by investigating the most general set of memory operations allowed in each theory, then establishing bounds on their capabilities.
The violations of classical bounds we have found showcase the importance of investigating classical memory effects in terms of manifestly classical descriptions, instead of relying on quantum-inspired analogies, as otherwise unexpected quantum effects may persist. As we have shown, the common presumption that EB channels act as a classical resource carries additional assumptions, which could be rather strong; see discussion on App. B. Furthermore, our results emphasize that, in the context of temporal correlations, it is essential to look beyond qubits and two-time scenarios, as otherwise important distinctions between classical and quantum systems are not apparent, as also recognized in Ref. [17]; see also App. D.
Our findings inspire several future research directions. Understanding the minimal requirements for an EB channel to truly act as a classical memory could offer insights into the origins of the quantum advantage in the context of temporal correlations, allowing a more definitive understanding of classical and quantum memory effects in non-Markovian processes; see also App. B. Furthermore, since the quantum advantage seen in this work stems from the non-orthogonality of the states $\sigma_{i}$, an investigation of the quantum upper bounds in the presence of EB channels could lead to new insights into the optimal encoding of classical information within quantum systems for various sequential tasks, with potential applications to, e.g., quantum random-access codes [33, 34, 35, 36] or quantum compression of classical predictive models for stochastic processes [37, 38, 39].
Naturally, there are many distinct features one could wish from a “quantum memory”, e.g., preserving coherence or entanglement [1, 9, 10], or enabling non-classical multi-time statistics [7, 17]. If the quantum memory is required to preserve entanglement, then EB channels are obviously useless as a quantum resource. However, quantum effects are not restricted to spatial correlations, and in considering the temporal character inherent to any information processing task, we are required to also consider the fundamental differences between classical and quantum temporal correlations. Our results indicate a more careful and detailed investigation of quantum memory resources is warranted, such that their precise origin and advantage over classical memories can be properly understood, quantified, and successfully applied in emergent quantum technologies.
Acknowledgements.— We thank Simon Milz and Philip Taranto for valuable discussions. This work is supported by the Austrian Science Fund (FWF) through projects ZK 3 (Zukunftskolleg) and F7113 (BeyondC). This work is supported by the Ministry of Science and Technology, Taiwan, (Grants No. MOST 111-2917-I-564-005, 112-2112-M-003-020-MY3), and Higher Education Sprout Project of National Taiwan Normal University (NTNU) and the Ministry of Education (MOE) in Taiwan.
Appendix A Direct correspondence between classical and quantum models for $m\leq d$
Using $\mathcal{A}$ for the set of outcomes, let $T=(T_{a})_{a\in\mathcal{A}}$ and $\{\left|i\right\rangle\}_{i=1}^{d}$ denote an arbitrary classical model and orthonormal basis, respectively. We consider a generic quantum instrument written in its Kraus decomposition
$\mathcal{I}_{a}(\rho)=\sum_{k=1}^{n_{a}}K_{ak}\rho K_{ak}^{\dagger},\quad\text{with}\quad\sum_{a\in\mathcal{A}}\sum_{k=1}^{n_{a}}K_{ak}^{\dagger}K_{ak}=\mathbbm{1},$ (9)
such that $\sum_{a}\mathcal{I}_{a}$ is trace preserving. We can define a Kraus operator for each non-zero $[T_{a}]_{ij}$, via the action $K_{aij}\left|i\right\rangle=\sqrt{[T_{a}]_{ij}}\left|j\right\rangle$, such that $\mathcal{I}$ acts diagonally in this basis. Now, by choosing $E_{i}=\sigma_{i}=\left|i\right\rangle\left\langle i\right|$, we have a completely dephasing channel $\mathcal{E}_{\text{CD}}$ in the same basis, $\mathcal{E}_{\text{CD}}(\rho):=\sum_{i=1}^{m}\left\langle i\right|\rho\left|i\right\rangle\left|i\right\rangle\left\langle i\right|$. Finally, by using the initial state $\rho_{0}=\sum_{i}\pi_{i}\left|i\right\rangle\left\langle i\right|$, we conclude
$\operatorname{tr}\left[\mathcal{I}_{a_{L}}\circ\mathcal{E}\circ\cdots\circ\mathcal{I}_{a_{1}}\circ\mathcal{E}(\rho_{0})\right]=\pi T_{a_{1}}\cdots T_{a_{L}}\eta.$ (10)
Together with the results from Eqs. 4 and 5, this establishes the correspondence between classical and quantum scenarios for $m\leq d$, and motivates the investigation of the $m>d$ case.
Appendix B Conditions for classicality
Since generic EB channels can violate bounds for classical memories, it is desirable to know the conditions under which non-classical memory effects are actually erased by their application, as often assumed, i.e., such that no choice of quantum instrument $\mathcal{I}$ can lead to non-classical temporal correlations.
A full investigation of these conditions is beyond the scope of this paper. Nevertheless, understanding some sufficient conditions offers insights on the strength of the implicit assumptions being made when applying EB channels as means of characterizing classical memory effects. We present two independent sufficient conditions: either the states $\{\sigma_{i}\}_{i=1}^{m}$ commute, and/or the POVM elements $\{E_{i}\}_{i=1}^{m}$ commute.
If $[\sigma_{i},\sigma_{j}]=0$, $\forall i,j$, then a common eigenbasis exists in which all operators can be written as $\sigma_{i}=\mathrm{diag}(s^{0}_{i},\dots,s^{d-1}_{i})$. Without loss of generality, we can express the channel in this basis as
$\begin{split}\mathcal{E}(\rho)=\sum_{i=1}^{m}\operatorname{tr}\left[E_{i}\rho\right]\sigma_{i}=\sum_{i=1}^{m}\sum_{\ell=0}^{d-1}\operatorname{tr}\left[E_{i}\rho\right]s_{i}^{\ell}\left|\ell\right\rangle\left\langle\ell\right|\\ =\sum_{\ell=0}^{d-1}\operatorname{tr}\left[\left(\sum_{i=1}^{m}s_{i}^{\ell}E_{i}\right)\rho\right]\left|\ell\right\rangle\left\langle\ell\right|=\sum_{\ell=0}^{d-1}\operatorname{tr}\left[F_{\ell}\rho\right]\left|\ell\right\rangle\left\langle\ell\right|,\end{split}$ (11)
where $F_{\ell}:=\sum_{i}s_{i}^{\ell}E_{i}$ and $\{F_{\ell}\}_{\ell=1}^{d}$ is a valid POVM. In fact, $s_{i}^{\ell}\geq 0$ since $\sigma_{i}\geq 0$ and $\sum_{\ell}s_{i}^{\ell}=1$ since $\operatorname{tr}\left[\sigma_{i}\right]=1$, which implies $F_{\ell}\geq 0$ and $𝟙\sum_{\ell}F_{\ell}=\sum_{i,\ell}s_{i}^{\ell}E_{i}=\sum_{i}E_{i}=\openone$. We thus have a new definition of the channel $\mathcal{E}$ which uses only $d$ outcomes and states. But from App. A, this implies there exists an equivalent classical model with $d$ states. The case for $[E_{i},E_{j}]=0$ is analogous.
Appendix C Numerical optimization
For $\bm{a}^{L}_{\text{ot}}=0\dots 01$ and $d=L-1$, the best classical bounds known are achieved by the so-called one-way models [6, 25]
$[T_{0}]_{i,i}=1/L,\quad[T_{0}]_{i,i+1}=[T_{1}]_{d,1}=1-1/L,$ (12)
with $p(\bm{a}^{L}_{\text{ot}}|T,d)=(1-\nicefrac{{1}}{{L}})^{L}$, which are conjectured to give the optimal probability for a given $d$ in any sub-deterministic classical scenario (i.e., $d<\operatorname{DC}(\bm{a})$). For $d=2,3$, and $4$, these upper bounds are known to be exact, and can be computed with semidefinite programming techniques [29].
In the quantum case, numerical optimization was performed via gradient descent techniques using the Adam algorithm [40], with an additional exponentially decreasing learning rate from $0.07$ to $10^{-12}$ over 50 000 iterations, and many trials with random initial values. We have found sufficient to optimize for rank-1 $E_{i},\sigma_{i}$, with no further assumptions. Each $\mathcal{I}_{a}$ was defined with a single Kraus operator, as no advantage was observed in using more. As the probability in Eq. 4 is convex for $\rho_{0}$, and we maximize for a single sequence, this initial state can be assumed pure, and w.l.o.g., set to $\rho_{0}=\left|0\right\rangle\left\langle 0\right|$. We have the following constrained optimization problem:
$\begin{split}\textbf{Find:}&\quad\max_{\mathcal{E},\mathcal{I}}p(\bm{a}|\mathcal{E},\mathcal{I},d)\\ \textbf{Such that:}&\quad E_{i},\sigma_{i}\geq 0\;\forall i,\\ &\quad\sum_{i}E_{i}=\mathbbm{1},\;\operatorname{tr}\left[\sigma_{i}\right]=1,\;\forall i,\\ &\quad\sum_{a}K_{a}^{\dagger}K_{a}=\mathbbm{1},\end{split}$ (13)
with $\mathcal{E}(\rho)=\sum_{i=1}^{m}\operatorname{tr}\left[\rho E_{i}\right]\sigma_{i}$ and $\mathcal{I}_{a}(\rho)=K_{a}\rho K_{a}^{\dagger}$. This can be relaxed into an unconstrained optimization by using arbitrary $d\times d$ complex matrices $A_{i},B_{i},C_{a}$, then defining $\sigma_{i}:=A_{i}^{\dagger}A_{i}/\mathrm{tr}[A_{i}^{\dagger}A_{i}]$ for the states. For the POVM, we first define $\tilde{B}_{i}:=B_{i}^{\dagger}B_{i}$ and compute $\lambda^{E}_{\text{max}}:=\max\text{eigs}(\sum_{i}\tilde{B}_{i})$, such that letting $E_{i}:=\tilde{B}_{i}/\lambda^{E}_{\text{max}}$ gives $\sum_{i}E_{i}\leq\mathbbm{1}$. For the instrument we compute $\lambda^{\mathcal{I}}_{\text{max}}=\max\text{eigs}(\sum_{a}C_{a}^{\dagger}C_{a})$ and define $K_{a}:=C_{a}/\sqrt{\lambda^{\mathcal{I}}_{\text{max}}}$, giving $\sum_{a}K_{a}^{\dagger}K_{a}\leq\mathbbm{1}$. In this way, the gradient of the objective function will naturally favor solutions satisfying $\sum_{i}E_{i}=\mathbbm{1}$ and $\sum_{a}K_{a}^{\dagger}K_{a}=\mathbbm{1}$. The above construction is general, but can be easily adapted for rank-1 $\sigma_{i}$ and $E_{i}$. The best quantum models found for $L=4$ and $5$ are included in App. D.
Appendix D Numerical examples violating the classical the bound
The following are the best quantum models found for $L=4$ and $5$ using gradient descent techniques, as described in App. C. These optimal values for a given dimension and sequence length required high accuracy numerical optimization to be found, and rely on several significant figures in order to adequately satisfy the constraints in Eq. 13. Despite considerable efforts, no closed-form description of these models was found. Since all optimal $E_{i}$ and $\sigma_{i}$ found were rank-1, we include the vectors such that $E_{i}=\left|e_{i}\right\rangle\left\langle e_{i}\right|$ and $\sigma_{i}=\left|\phi_{i}\right\rangle\left\langle\phi_{i}\right|$. We provide only 4 significant digits of precision, for simplicity, but convergence was obtained up to 10 significant digits. Both cases assume $\rho_{0}=\left|0\right\rangle\left\langle 0\right|$ and $\mathcal{I}_{a}(\rho)=K_{a}\rho K_{a}^{\dagger}$.
Model for $L=4$:
$\begin{split}&\left|e_{1}\right\rangle=\begin{bmatrix}1\\ 0\\ 0\end{bmatrix},\quad\left|e_{2}\right\rangle=\begin{bmatrix}0\\ -0.09692-0.41924i\\ 0.74404\end{bmatrix},\quad\left|e_{3}\right\rangle=\begin{bmatrix}0\\ 0.07184\,+0.30898i\\ 0.65475\end{bmatrix},\quad\left|e_{4}\right\rangle=\begin{bmatrix}0\\ -0.18847-0.82383i\\ -0.13307\end{bmatrix},\\ &\left|\phi_{1}\right\rangle=\begin{bmatrix}0\\ -0.08293-0.35949i\\ 0.92946\end{bmatrix},\quad\left|\phi_{2}\right\rangle=\begin{bmatrix}0\\ 0.09971\,+0.4288i\\ 0.89788\end{bmatrix},\quad\left|\phi_{3}\right\rangle=\begin{bmatrix}0\\ 0.96462\\ 0.05913\,-0.25695i\end{bmatrix},\quad\left|\phi_{4}\right\rangle=\ \begin{bmatrix}1\\ 0\\ 0\end{bmatrix},\\ &K_{0}=\mathrm{diag}(0,1,1),\qquad K_{1}=\mathrm{diag}(1,0,0).\end{split}$ (14)
Model for $L=5$:
$\begin{split}&\left|e_{1}\right\rangle=\begin{bmatrix}1\\ 0\\ 0\\ 0\end{bmatrix},\quad\left|e_{2}\right\rangle=\begin{bmatrix}0\\ \ -0.18168-0.03287i\\ -0.33582-0.01918i\\ 0.83859\end{bmatrix},\quad\left|e_{3}\right\rangle=\ \begin{bmatrix}0\\ 0.06342\,+0.62645i\\ -0.36075+0.0699i\\ -0.32475\end{bmatrix},\quad\ \left|e_{4}\right\rangle=\begin{bmatrix}0\\ -0.40607-0.42989i\\ -0.28345-0.24945i\\ -0.4251\end{bmatrix}\ ,\quad\left|e_{5}\right\rangle=\begin{bmatrix}0\\ -0.3906+0.2592i\\ 0.78055\\ 0.05633\,+0.08616i\ \end{bmatrix},\\ &\left|\phi_{1}\right\rangle=\begin{bmatrix}0\\ -0.25627-0.10917i\\ -0.29951-0.06935i\\ 0.90988\ \end{bmatrix},\quad\left|\phi_{2}\right\rangle=\begin{bmatrix}0\\ 0.78968\\ 0.17386\,+0.37575i\\ \ -0.02481+0.45209i\end{bmatrix},\quad\left|\phi_{3}\right\rangle=\begin{bmatrix}0\\ 0.73235\\ 0.46844\,\ -0.13164i\\ 0.35015\,-0.32293i\end{bmatrix},\quad\left|\phi_{4}\right\rangle=\begin{bmatrix}0\\ \ -0.34778+0.40375i\\ 0.82555\\ 0.14552\,+0.11544i\end{bmatrix},\quad\left|\phi_{5}\right\rangle=\ \begin{bmatrix}1\\ 0\\ 0\\ 0\end{bmatrix},\\ &K_{0}=\mathrm{diag}(0,1,1,1),\qquad K_{1}=\mathrm{diag}(1,0,0,0).\end{split}$ (15)
For $L=3$ and $d=2$, we were unable to find any violation of the classical bound, strongly suggesting EB-based temporal non-classicality only emerges beyond two-level systems. Furthermore, since $\operatorname{DC}(\bm{a})\leq L$ for any sequence of length $L$, either a qubit or a bit are sufficient to produce any correlations on two time steps. It is therefore essential to also look beyond two-time steps in order to reveal fundamental differences between classical and quantum systems in this sequential measurement scenario.